# Deep Optimal Transport for Domain Adaptation on SPD Manifolds

### v4

@@ II-C Joint Distribution Optimal Transport
Assuming that the transfer between source and target distributions can be estimated using optimal transport, Joint Distribution Optimal Transport (JDOT) [8] extends the OT-DA framework by jointly optimizing both the optimal coupling and the prediction function $f$ for the two distributions, i.e., $\min_{\gamma\in\prod(\mu,\nu)}\big{\langle}\gamma,c^{\prime}(x,y;\,\bar{x},f(\bar{x}))\big{\rangle}_{\mathcal{F}}$, where $y$ is the label set for samples $x$, $f$ is the prediction function, and the joint cost function is a linear combination of distance function and cross-entropy loss, i.e., $c^{\prime}(x,y;\bar{x},f(\bar{x})):=\alpha_{1}d_{\ell^{2}}(x,\bar{x})+\alpha_{2}\mathcal{L}(y,f(\bar{x}))$. Moreover, DeepJDOT employs neural networks to map raw data into feature space and thus extends JDOT by setting the new joint objective function $\mathcal{L}:=\mathcal{L}(y,f(g(x)))+\big{\langle}\gamma,c^{\prime}\big{(}g(x),y;\,g(\bar{x}),f(g(\bar{x})\big{)}\big{\rangle}_{\mathcal{F}}$, where $g$ represents the neural networks [9].
II-D SPD Matrix-Valued Motor Imagery Classifier
Extensive research has explored various classes of Riemannian metrics on SPD manifolds for various applications. These explorations have led to the introduction of numerous statistical methods to address computational issues on SPD manifolds, drawing inspiration from manifold-valued neuroimaging data [19, 20]. During the feature engineering era, Barachant et al. utilized the Riemannian distance between EEG spatial covariance matrices for motor imagery classification, resulting in a significant impact on various brain-computer interface applications [21, 22, 23].
In the deep learning era, Ju et al. employed geometric deep learning on SPD manifolds to extract discriminative information from the time-space-frequency domain for motor imagery classification, ultimately enhancing classification performance [24, 25, 26, 27, 28]. The neural network layers utilized in these works are derived from SPDNet [29], which is specifically designed to process SPD matrices while preserving their SPD properties throughout the layers during non-linear learning. These layers include:
@@ II-D SPD Matrix-Valued Motor Imagery Classifier
  • Bi-Map Layer: This layer transforms the covariance matrix $S$ by applying the Bi-Map transformation $WSW^{T}$, where the transformation matrix $W$ must have full row rank.
@@
  • LOG Layer: This layer maps elements on SPD manifolds to their tangent space using $U\log{(\Sigma)}U^{T}$, where $S=U\Sigma U^{T}$. The logarithm function applied to a diagonal matrix takes the logarithm of each diagonal element individually.
III Methodology
@@ III-A OT-DA Framework on ($\mathcal{S}_{++}$, $g^{LEM}$)
Without loss of generality, consider a point $p\in\mathcal{M}^{\circ}$. There exists an open neighborhood $\mathcal{N}$ of $p$ and $\epsilon>0$ such that the exponential map $\exp_{p}$ maps the ball $\mathcal{B}(0,\epsilon)\subset T_{p}\mathcal{M}$ diffeomorphically onto $\mathcal{N}$ [30, Corollary 5.5.2.]. Let $c(s)$ be a geodesic on $\mathcal{N}$, parameterized with constant speed, with $c(0):=q\in\mathcal{N}$ and $c(1):=p$ with $\dot{c}(0)=\log_{q}(p)$. By reversing the direction of the geodesic $c(1-s)$, we have $\dot{c}(1)=-\log_{p}(q)$. We proceed by linearizing the exponential maps around $0\in T_{p}\mathcal{M}$ and $\dot{c}(0)\in T_{q}\mathcal{M}$ and derive the derivative of a function $f$ as follows:
@@
where the differential $(D\exp_{p})_{q}(\cdot):T_{q}(T_{p}\mathcal{M})\longmapsto(T_{p}\mathcal{M})$ is the differential of the exponential map at $q$ with respect to $p$. The fourth equality follows from $(D\exp_{p})_{0}=I$ on $T_{p}\mathcal{M}$, and the fifth equality follows from Gauss’ lemma [31, Section 3.3.5.]. Hence, by the definition of the gradient on Riemannian manifolds, we have $df(p)(v)=g_{p}\big{(}\nabla f(x),v\big{)}$, which implies that $\nabla f(p)=\dot{c}(1)=-\log_{p}(q)$. ∎
@@
where $\prod(\mu,\nu):=\{\gamma\in\mathbb{R}_{\geq 0}^{N\times M}\big{|}\gamma\mathbbm{1}_{N}=\frac{1}{N},\text{ and }\gamma^{\top}\mathbbm{1}_{M}=\frac{1}{M}\}$. The pseudocode for the computation of the transport plan is presented in Algorithm 1.
Input  : Source samples $\{S_{i}\}_{i=1}^{N}\sim\mu$, target samples $\{\bar{S}_{j}\}_{j=1}^{M}\sim\nu$.
Output: Transport plan $\gamma$, and new coordinates for target samples $\{\bar{S}^{\dagger}_{j}\}_{j=1}^{M}$.
/* 1. Compute ground metric matrix for OT */
@@
/* 2. Compute transport plan $\gamma$ using the earth movers distance-based OT algorithm. */
          $\gamma^{\dagger}\leftarrow OT^{EMD}(N,M,D(i,j))$.
/* 3. Compute new coordinates using transport plan on ($\mathcal{S}_{++}$, $g^{LEM}$). */
          $\{\bar{S}^{\dagger}_{j}\}_{j=1}^{M}\leftarrow\exp{\big{(}\gamma^{\dagger}[j,:]\cdot\{\log(S_{i})\}_{i=1}^{N}\big{)}}$.
@@
On ($\mathcal{S}_{++}$, $g^{LEM}$), consider the source measure $\mu$ and target measure $\nu\in\mathbb{P}(\mathcal{S}_{++})$, which are only accessible through finite samples $\{S_{i}\}_{i=1}^{N}\sim\mu$ and $\{\bar{S}_{j}\}_{j=1}^{M}\sim\nu$. Each sample contributes to the empirical distribution with weights $\frac{1}{N}$ and $\frac{1}{M}$, respectively. Assuming the existence of a linear transformation between the feature space distributions of the source and target domains, it can be estimated using the discrete Monge-Kantorovich Problem 2. Therefore, the values of the continuous optimal transport $T(S)$ lie within the set $\big{\{}\exp{\big{(}\gamma^{\dagger}[j,:]\cdot{\log(S_{i})}_{i=1}^{N}\big{)}}\big{\}}_{j=1}^{M}$.
Theorem 3 encompasses an extension of [6, Theorem 3.1] pertaining to ($\mathcal{S}_{++}$, $g^{LEM}$) within the context of the OT-DA framework. Significantly, our framework is designed to incorporate any supervised neural networks-based transformation, distinguishing it from the conventional use of affine transformations, and employs the squared Riemannian distance. In Appendix VI-A, we illuminate the connection between our proposed approach and another similar OT-DA framework on SPD manifolds proposed in [10] that utilizes the squared $\ell_{2}$ Euclidean distance in its cost function.
Figure 1: Illustration of Deep Optimal Transport: Multi-channel signals are first transformed into covariance matrices, situating these matrices on the space of ($\mathcal{S}_{++}$, $g^{LEM}$). The SPD matrix-valued data is processed through the Bi-Map, ReEig, and LOG layers, moving it to the tangent space at the identity. The loss function takes into account the cross-entropy loss, marginal distribution adaptation, and conditional distribution adaptation.
III-B Deep Optimal Transport
We propose a novel geometric deep learning approach, Deep Optimal Transport (DOT) on ($\mathcal{S}_{++}$, $g^{LEM}$), to address the joint distribution adaptation problems involving both marginal distribution shifts and conditional distribution shifts. We assume samples from the source and target distributions can be restored via OT. The proposed architecture is depicted in Figure 1.
@@ III-B Deep Optimal Transport
Conditional Distribution Adaptation (CDA)

### v5

@@ 2.3 Joint Distribution Optimal Transport
where $y$ denotes the label set for samples $x$, $f$ is the prediction function, and the joint cost function combines a distance metric with a cross-entropy loss: $c(x,y;\bar{x},f(\bar{x})):=\alpha_{1}d_{\ell^{2}}(x,\bar{x})+\alpha_{2}\mathcal{L}(y,f(\bar{x}))$. DeepJDOT further enhances this framework by employing neural networks to embed raw data into a feature space. Its joint objective function becomes: $\mathcal{L}_{total}:=\mathcal{L}(y,f(g(x)))+\big{\langle}\gamma,c\big{(}g(x),y;\,g(\bar{x}),f(g(\bar{x})\big{)}\big{\rangle}_{\mathcal{F}}$, where $g$ represents the neural network encoder [14].
2.4 EEG-Based Motor Imagery Classifier on SPD Manifolds
Extensive research has explored various classes of Riemannian metrics on SPD manifolds, with applications spanning neuroimaging and machine learning. These efforts have led to the development of manifold-valued statistical methods aimed at addressing computational challenges inherent to SPD manifolds—particularly in analyzing neuroimaging data with intrinsic geometric structures [18, 19]. In the feature engineering era, Barachant et al. pioneered the use of Riemannian distances between EEG spatial covariance matrices for motor imagery classification, marking a major advancement in brain-computer interface research. Their method achieved strong performance in both multiclass classification and real-time applications [10, 20, 21]. Building on this foundation, Ju et al. advanced geometric deep learning methods on SPD manifolds in the deep learning era, developing architectures that extract discriminative spatio-temporal-frequency features [22, 23, 24, 25, 26]. The neural network-based architecture in these works originated from SPDNet [27], a foundational architecture designed to process SPD matrices while preserving their manifold structure through the following layers:
@@ 2.4 EEG-Based Motor Imagery Classifier on SPD Manifolds
  • BiMap Layer: This layer transforms the covariance matrix $S$ by applying the BiMap transformation $WSW^{T}$, where the transformation matrix $W$ must have full row rank.
@@
  • LogEig Layer: This layer maps elements on SPD manifolds to their tangent space using $U\log{(\Sigma)}U^{T}$, where $S=U\Sigma U^{T}$. The logarithm function applied to a diagonal matrix takes the logarithm of each diagonal element individually.
3 Methodology
@@ 3.1 OT-DA Framework on ($\mathcal{S}_{++}$, $g^{LEM}$)
Without loss of generality, consider a point $p\in\mathcal{M}^{\circ}$. There exists an open neighborhood $\mathcal{N}$ of $p$ and $\epsilon>0$ such that the exponential map $\exp_{p}$ maps the ball $\mathcal{B}(0,\epsilon)\subset T_{p}\mathcal{M}$ diffeomorphically onto $\mathcal{N}$ [28, Corollary 5.5.2.]. Let $c(s)$ be a geodesic on $\mathcal{N}$, parameterized with constant speed, with $c(0):=q\in\mathcal{N}$ and $c(1):=p$ with $\dot{c}(0)=\log_{q}(p)$. By reversing the direction of the geodesic $c(1-s)$, we have $\dot{c}(1)=-\log_{p}(q)$. We proceed by linearizing the exponential maps around $0\in T_{p}\mathcal{M}$ and $\dot{c}(0)\in T_{q}\mathcal{M}$ and derive the derivative of a function $f$ as follows:
@@
where the differential $(D\exp_{p})_{q}(\cdot):T_{q}(T_{p}\mathcal{M})\longmapsto(T_{p}\mathcal{M})$ is the differential of the exponential map at $q$ with respect to $p$. The fourth equality follows from $(D\exp_{p})_{0}=I$ on $T_{p}\mathcal{M}$, and the fifth equality follows from Gauss’ lemma [29, Section 3.3.5.]. Hence, by the definition of the gradient on Riemannian manifolds, we have $df(p)(v)=g_{p}\big{(}\nabla f(x),v\big{)}$, which implies that $\nabla f(p)=\dot{c}(1)=-\log_{p}(q)$. ∎
@@
where transportation plan is as follows,
$\prod(\mu,\nu):=\{\gamma\in\mathbb{R}_{\geq 0}^{N\times M}\big{|}\gamma\mathbbm{1}_{N}=\frac{1}{N},\text{ and }\gamma^{\top}\mathbbm{1}_{M}=\frac{1}{M}\}.$
The pseudocode for the computation of the transport plan is presented in Algorithm 1.
Input  : Source samples $\{S_{i}\}_{i=1}^{N}\sim\mu$ and target samples $\{\bar{S}_{j}\}_{j=1}^{M}\sim\nu$.
Output: Transport plan $\gamma$ and new coordinates for target samples $\{\bar{S}^{\dagger}_{j}\}_{j=1}^{M}$.
1. Compute ground metric matrix for OT as follows:
@@
2. Compute transport plan $\gamma$ using the earth movers distance-based OT algorithm as follows:                       $\gamma^{\dagger}\leftarrow OT^{EMD}(N,M,D(i,j))$.
3. Compute new coordinates using transport plan on ($\mathcal{S}_{++}$, $g^{LEM}$) as follows:                       $\{\bar{S}^{\dagger}_{j}\}_{j=1}^{M}\leftarrow\exp{\big{(}\gamma^{\dagger}[j,:]\cdot\{\log(S_{i})\}_{i=1}^{N}\big{)}}$.
@@
On ($\mathcal{S}_{++}$, $g^{LEM}$), consider the source measure $\mu$ and target measure $\nu\in\mathbb{P}(\mathcal{S}_{++})$, which are only accessible through finite samples $\{S_{i}\}_{i=1}^{N}\sim\mu$ and $\{\bar{S}_{j}\}_{j=1}^{M}\sim\nu$. Each sample contributes to the empirical distribution with weights $\frac{1}{N}$ and $\frac{1}{M}$, respectively. Assuming the existence of a linear transformation between the feature space distributions of the source and target domains, it can be estimated using the discrete Monge-Kantorovich Problem 2. Therefore, the values of the continuous optimal transport $T(S)$ lie in the set $\big{\{}\exp{\big{(}\gamma^{\dagger}[j,:]\cdot{\log(S_{i})}_{i=1}^{N}\big{)}}\big{\}}_{j=1}^{M}$.
Remark.
1). Theorem 3 extends [5, Theorem 3.1] to the OT-DA framework on SPD manifolds equipped with the Log-Euclidean metric. Unlike conventional OT-DA methods restricted to affine transformations, our framework uniquely incorporates arbitrary supervised neural network-based mappings while minimizing transportation costs via the squared Riemannian distance $d_{g^{LEM}}^{2}(\cdot,\cdot)$.
2). In Appendix 6.1, we contrast our approach with the SPD manifold OT-DA framework in [11], which employs the squared Euclidean distance as its cost function. This comparison highlights the geometric advantages of our Riemannian distance formulation over Euclidean approximations.
Figure 1: Illustration of Deep Optimal Transport: Multi-channel signals are first transformed into covariance matrices, situating these matrices on the space of ($\mathcal{S}_{++}$, $g^{LEM}$). The SPD matrix-valued data is processed through the BiMap, ReEig, and LogEig layers, moving it to the tangent space at the identity. The loss function takes into account the cross-entropy loss, marginal distribution adaptation, and conditional distribution adaptation.
3.2 Deep Optimal Transport
We propose Deep Optimal Transport (DOT), a SPDNet-based architecture on ($\mathcal{S}_{++}$, $g^{LEM}$), to address joint distribution adaptation under both marginal and conditional shifts. Unlike traditional OT methods, DOT leverages a deep encoder to learn nonlinear push-forward mappings that align source and target distributions by minimizing a composite OT loss. The key assumption is that domain discrepancies can be bridged through transport plans parameterized by neural networks. As illustrated in Figure 1, the architecture integrates:
  1. i.
    Nonlinear Manifold Mapping $\varphi:X_{S},X_{T}\longmapsto Z$: Using multiple BiMap layers and ReEig layers from SPDNet as transformations on SPD manifolds.
  2. ii.
    Projection: Applying the LogEig layer from SPDNet to flatten the SPD manifold-valued data onto the Euclidean tangent space at identity.
  3. iii.
    Joint OT-Driven Distribution Alignment: Simultaneously minimizing the OT costs for marginal and conditional distributions as follows:
@@ 3.2 Deep Optimal Transport
    The primary objective of MDA is to align the marginal distributions $\mathbb{P}_{S}\big{(}\varphi(X_{S})\big{)}$ and $\mathbb{P}_{T}\big{(}\varphi(X_{T})\big{)}$ on the Riemannian manifold ($\mathcal{S}_{++}$, $g^{LEM}$). We quantify their statistical discrepancy through the Riemannian distance between their Log-Euclidean Fréchet means. Specifically, for batches of source and target samples $\mathcal{B}_{S}$ and $\mathcal{B}_{T}$, the MDA loss is defined as the Frobenius norm of the difference between their logarithmic mean matrices:
    $\mathcal{L}_{MDA}:=||\log\big{(}\overline{w}(\mathcal{B}_{S})\big{)}-\log\big{(}\overline{w}(\mathcal{B}_{T})\big{)}||_{\mathcal{F}}.$
    Conditional Distribution Adaptation (CDA)
    CDA aims to match the class-conditional distributions $\mathbb{Q}_{S}(\varphi(Y_{S})\big{|}\varphi(X_{S}))$ and $\mathbb{Q}_{T}(\varphi(Y_{T})\big{|}\varphi(X_{T}))$ on the same manifold ($\mathcal{S}_{++}$, $g^{LEM}$), contrasting with traditional posterior probability adaptation. In unsupervised settings, pseudo-labels $\hat{y}$ for target domain samples (generated by a baseline algorithm) are utilized to construct class-specific sample batches. For each class $\ell\in\{1,....,L\}$, we define $S^{\ell}:=\{x_{S}|\hat{y}_{S}={\ell}\}$ and $T^{\ell}:=\{x_{T}|\hat{y}_{T}=\ell\}$, then compute the Riemannian distance between their Fréchet means. The CDA loss is formulated as:
    $\mathcal{L}_{CDA}:=||\log\big{(}\overline{w}(\mathcal{B}_{S^{\ell}})\big{)}-\log\big{(}\overline{w}(\mathcal{B}_{T^{\ell}})\big{)}||_{\mathcal{F}}.$
    Overall Objective Function
    The overall objective function of the DOT framework combines the cross-entropy loss $\mathcal{L}_{CE}$ with joint distribution adaptation terms, formulated as:
    $\mathcal{L}_{DOT}:=\alpha_{1}\mathcal{L}_{CE}+\alpha_{2}\mathcal{L}_{MDA}^{2}+\alpha_{3}\mathcal{L}_{CDA}^{2},$
    where $\alpha_{1},\alpha_{2}$, and $\alpha_{3}\geq 0$ are weighting coefficients governing the trade-off between classification accuracy and distribution alignment.
Remark.
1). Based on Brenier’s theorem and related theoretical results [30], both $\mathcal{L}_{MDA}$ and $\mathcal{L}_{CDA}$ adopt squared-distance formulations.
@@
3). By substituting Equation 1 into $\mathcal{L}_{MDA}$ and $\mathcal{L}_{CDA}$, we obtain:
$\mathcal{L}_{MDA}=||\sum_{S_{i}\in\mathcal{B}_{S}}\frac{\log{(S_{i})}}{|\mathcal{B}_{S}|}-\sum_{S_{j}\in\mathcal{B}_{T}}\frac{\log{(S_{j})}}{|\mathcal{B}_{T}|}||_{\mathcal{F}};\hskip 10.00002pt\mathcal{L}_{CDA}=||\sum_{{S^{\ell}_{i}}\in\mathcal{B}_{S^{l}}}\frac{\log{({S^{\ell}_{i}})}}{|\mathcal{B}_{S^{\ell}}|}-\sum_{{S^{\ell}_{j}}\in\mathcal{B}_{T^{\ell}}}\frac{\log{({S^{\ell}_{j}})}}{|\mathcal{B}_{T^{\ell}}|}||_{\mathcal{F}}.$
These simplified forms have dual interpretations:
  • Empirical Maximum Mean Discrepancy (MMD): Let $n=|\mathcal{B}_{S}|$, $m=|\mathcal{B}_{T}|$, $\varphi{(x)}:=\log{(x)}$, and $\mathcal{H}:=\mathbb{R}^{d}$. The loss becomes $\mathcal{L}_{MMD}=||\sum_{i=1}^{n}\varphi{(S_{i})}/n-\sum_{j=1}^{m}\varphi{(S_{i})}/m||_{\mathcal{H}}$.
  • Correlation Alignment (CORAL): As covariance matrices of multi-channel signals, the alignment of second-order statistics minimizes domain drift [31, 32].
4). Multi-Source Adaptation: For multi-source scenarios, we extend the loss by integrating pairwise source-target adaptations on ($\mathcal{S}_{++}$, $g^{LEM}$):
$\mathcal{L}:=\alpha_{1}^{i}\mathcal{L}_{CE}+\sum_{i}\Big{(}\alpha_{2}^{i}\mathcal{L}_{MDA,i}+\alpha_{3}^{i}\mathcal{L}_{CDA,i}\Big{)},$
where $i^{th}$ indexes source domains. This aligns with multi-marginal optimal transport theory on manifolds [33] and multi-source joint transport [34].
5). DOT Framework: DOT does not explicitly solve an OT problem. Instead, it leverages transport costs as a guiding principle for neural network training. The trained network maps source and target domains onto a submanifold of the SPD manifold, effectively implementing a transfer learning subspace method.
6). Adaptation Strategy: Unlike DeepJDOT, which computes transport costs for all points and requires alternating parameter updates due to its deterministic transport plan, DOT calculates costs only for centroids and achieves conditional marginal adaptation without alternating updates.
7). Perspective from Bures-Wasserstein Geometry: Bures-Wasserstein Geometry is a geometric framework on SPD manifolds, where the 2-Wasserstein distance between centered Gaussian distributions induces the Bures metric, governing optimal transport and barycenter computations for covariance matrices [35]. Specifically, for centered Gaussian distributions $\mathcal{N}(0,C_{s})$ and $\mathcal{N}(0,C_{t})$, the 2-Wasserstein distance reduces to the Bures metric on the SPD manifold:
$W_{2}^{2}\big{(}\mathcal{N}(0,C_{s}),\mathcal{N}(0,C_{t})\big{)}=\text{Tr}\left(C_{s}+C_{t}-2\left(C_{s}^{1/2}C_{t}C_{s}^{1/2}\right)^{1/2}\right).$
The barycenter of Gaussians $\{\mathcal{N}(0,C_{i})\}_{i=1}^{n}$ under $W_{2}$ is defined via their covariances:
$\bar{C}=\arg\min_{C\succ 0}\sum_{i=1}^{n}W_{2}^{2}\big{(}\mathcal{N}(0,C),\mathcal{N}(0,C_{i})\big{)},$
with efficient computation via fixed-point iterations [36, 37]. When the deep encoder $\varphi$ is identity (i.e., $C_{i}=x_{i}x_{i}^{T}$) and data follows centered Gaussians with commuting covariances ($C_{s}C_{t}=C_{t}C_{s}$), $\mathcal{L}_{MDA}$ under $d_{\text{LEM}}$ recovers the Bures metric. This is because the logarithmic transformation under $g^{LEM}$ aligns with the geometric mean in Bures for commuting matrices. However, the DOT framework extends beyond this idealized scenario in two ways:
  1. i.
    The deep encoder $\varphi$ enables nonlinear manifold mapping, breaking the Gaussian assumption;
  2. ii.
    The composite loss $\mathcal{L}_{\text{CDA}}$ jointly aligns conditionals, which is nontrivial under pure Wasserstein barycenter frameworks.
4 EXPERIMENTS
4.1 Deep Optimal Transport-Based Motor Imagery Classifier
We model EEG spatial covariance matrices on the Riemannian manifold ($\mathcal{S}_{++}$, $g^{LEM}$) and employ DOT to address statistical distribution shifts during feedback phases. Motor imagery classification leverages event-related desynchronization/synchronization effects [38], which identify discriminative EEG frequency components in specific bands. The BiMap layer projects source and target features into a shared latent space where $X_{S}\approx X_{T}$, enabling SPD manifold modeling with the Log-Euclidean metric. This layer induces a transitive group action on the manifold, allowing smooth transitions between any two SPD points (Fig. 2 illustrates the DOT-based classifier architecture).
Formally, an EEG trial is represented by a matrix $X\in\mathbb{R}^{n_{C}\times n_{T}}$, where $n_{C}$ denotes the number of channels and $n_{T}$ the number of time points. To capture localized spatio-spectral dynamics, we compute spatial covariance matrices $S(\Delta f,\Delta t)\in\mathcal{S}_{++}^{n_{C}}$, where each matrix characterizes the signal within a specific frequency band $\Delta f$ and time interval $\Delta t$. Let $\mathcal{F}$ and $\mathcal{T}$ denote the overall frequency and time domains, respectively. A segmentation over these domains is defined as
$\text{segment}(\mathcal{F},\mathcal{T}):=\big{\{}(\Delta f\times\Delta t)\,\big{|}\,\Delta f\subseteq\mathcal{F},\Delta t\subseteq\mathcal{T}\big{\}}.$
To ensure practical convergence in real-world scenarios, the segmentation must satisfy the following conditions:
Assumption 1 (Necessary Optimal Transport Condition).
There exist linear transformations between the feature space distributions on the source and target domains for each segment of $\text{segment}(\mathcal{F},\mathcal{T})$, which can be estimated using the Kantorovich Problem 2.
Assumption 2 (Uniform Spectral Prior).